If two dies are toced together,what is the probability of having equal outcomes

1/6
5/36
1
6
1
Expert's answer
2015-03-03T06:19:29-0500
All elementary outcomes can be represented in the following table:
(1,1) (2,1) (3,1) (4,1) (5,1) (6,1)
(1,2) (2,2) (3,2) (4,2) (5,2) (6,2)
(1,3) (2,3) (3,3) (4,3) (5,3) (6,3)
(1,4) (2,4) (3,4) (4,4) (5,4) (6,4)
(1,5) (2,5) (3,5) (4,5) (5,5) (6,5)
(1,6) (2,6) (3,6) (4,6) (5,6) (6,6)
The number of elementary outcomes equal to 36. Since the throws are independent from each other, the elementary outcomes - are equally.
We denote the random event A - equal points. Favorable outcomes of this event: (1,1), (2,2), (3,3), (4,4), (5,5), (6,6). Only 6 outcomes. Hence, the probability that the sum of the two dice rolled eight points, equal to:
P(A)=6/36=1/6